# Partial-Label Learning with a Reject Option

### v3

@@ 3.2 Dealing with Uncertainty
Recently, much attention has been given to quantifying the uncertainty associated with predictions by machine-learning models. These include Bayesian methods (Kingma & Welling, 2014; Kendall & Gal, 2017) and ensembles (Lakshminarayanan et al., 2017; Wimmer et al., 2023). Notably, a probability measure $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{M}}_{1}^{+}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ cannot distinguish uncertainty arising from (i) and (ii) (Hüllermeier & Waegeman, 2021). To mitigate this limitation, we employ Dempster-Shafer theory (Dempster, 1967; Shafer, 1986). We argue that DST is a natural fit for PLL: With DST, we can assign probability mass to sets of candidates without specifying the probabilities of individual labels; incorrect labels do not obtain any probability mass. In contrast, existing PLL approaches (Hüllermeier & Beringer, 2005; Cour et al., 2011; Liu & Dietterich, 2012; Zhang & Yu, 2015; Ni et al., 2021; Xu et al., 2023) initially attribute some probability to each label candidate and subsequently refine them. By doing so, most probability mass is initially allocated to labels that are certainly incorrect, as only one candidate is the true label. Several methods already leverage DST in supervised learning (Mandler & Schümann, 1988; Denoeux, 1995; Tabassian et al., 2012; Sensoy et al., 2018; Denoeux, 2019; Tong et al., 2021). We consider the nearest neighbor approach by Denoeux (1995) to be most closely related to our approach. Basic probability assignments are constructed from the nearest neighbors of an instance. Then, Dempster’s rule is used to combine them into a single bpa. Their analysis is, however, not transferable to our case because they only have singletons or the full label space as focal sets making set intersections in the combination rule easy to handle. We examine a more general setting since probability mass is allocated to arbitrary subsets.
In the PLL context, we consider (Hüllermeier & Beringer, 2005; Zhang & Yu, 2015; Zhang et al., 2016; Xu et al., 2019; Wang et al., 2019) as most closely related to our proposed method as these approaches consider an instance’s neighborhood to predict its class label. To the best of our knowledge, we are the first to propose an extension of the $k$-NN classifier that keeps uncertainty from (i) and (ii) separate by leveraging Dempster-Shafer theory in the PLL context.
@@ 4 Our Method: DST-PLL
Algorithm 1 outlines our proposed method Dst-Pll, which we summarize in the following. We denote by $\operatorname{NN}_{k}(\tilde{x})\subseteq\operatorname{\mathcal{X}}\times 2^{\operatorname{\mathcal{Y}}}$ the set of the $k$-nearest neighbors of instance $\tilde{x}$ with their associated candidate labels. To predict the class label of an instance $\tilde{x}$, the algorithm first transforms information from $\tilde{x}$’s neighbors $\operatorname{NN}_{k}(\tilde{x})$ into bpas $\operatorname{m}_{i}$ (Lines 3–7), collects these into evidence set $\mathcal{E}$ (Line 8), and combines the bpas into $\operatorname{\tilde{m}}$ using Yager’s rule (Yager, 1987a, b) (Line 10). Section 4.1 elaborates on these steps and Section 4.2 details how we extract a prediction (Line 11). The latter also shows that the resulting classification rule is risk-consistent. Section 4.3 elaborates on how we extract our reject option from $\operatorname{\tilde{m}}$ (Line 12). We analyze our algorithm’s runtime in Section 4.4.
@@
4.1 Evidence from Partially-labeled Instances
Basic probability assignments. Following the standard assumption that neighboring instances in feature space are also close in label space, we combine the evidence from the $k$ nearest neighbors $(x_{i},s_{i})\in\operatorname{NN}_{k}(\tilde{x})$ of a given instance $\tilde{x}\in\operatorname{\mathcal{X}}$ with its candidate labels $\tilde{s}\subseteq\operatorname{\mathcal{Y}}$ ($\tilde{s}=\operatorname{\mathcal{Y}}$ if $\tilde{x}$ is a test instance). For fixed $i\in[k]$, the candidate labels $s_{i}$ do not provide any valuable information if they support all ($\tilde{s}\subseteq s_{i}$) or none ($\tilde{s}\cap s_{i}=\emptyset$) of the labels in $\tilde{s}$; we use a bpa of $\operatorname{m}_{i}(\tilde{s})=1$ (Line 4). We set $\operatorname{m}_{i}(A)=1/2$ if $A=\tilde{s}$ or $A=\tilde{s}\cap s_{i}$, else $\operatorname{m}_{i}(A)=0$ (Line 6), where $1/2$ equally weights evidence. We later elaborate further on this. Note that we make the common assumption that the true label of instance $\tilde{x}$ is always in $\tilde{s}$ (Cour et al., 2011; Liu & Dietterich, 2012; Lv et al., 2020; Ni et al., 2021). While our definition of the $\operatorname{m}_{i}$-s is similar to (Denoeux, 1995), we target a more general setting. Our focal sets can be arbitrary subsets instead of only singletons or the full label set.
@@ 4.1 Evidence from Partially-labeled Instances
Evidence combination. Given the set =$\mathcal{E}=\{\operatorname{m}_{i}\mid i\in[k]\}$, we combine all $\operatorname{m}_{i}$-s using Yager’s rule (Yager, 1987a, b). Dempster’s original rule (Dempster, 1967) enforces $\operatorname{\tilde{m}}(\emptyset)=0$ by normalization, which is criticized for its unintuitive results when facing high conflict (Zadeh, 1984). Instead, Yager’s rule first collects overlapping evidence in $\operatorname{q}:2^{\operatorname{\mathcal{Y}}}\rightarrow[0,1]$ and creates a valid bpa $\operatorname{\tilde{m}}:2^{\operatorname{\mathcal{Y}}}\rightarrow[0,1]$ by
@@ 4.2 Classification Rule and Consistency
After the combination into $\operatorname{\tilde{m}}$ by (4), we extract a prediction $g(\tilde{x})$ for instance $\tilde{x}$ (Line 11). We predict the class label with the highest probability mass $\arg\max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\{y\})$ if any has non-zero mass, or else randomly pick from the subset with the most mass $\arg\max_{A\subseteq\tilde{s}}\,\operatorname{\tilde{m}}(A)$. When tied, we use the subset with the smallest cardinality.
@@
Our classification rule yields a risk-consistent classifier, which we demonstrate in the following. As is common in the literature (Cour et al., 2011; Liu & Dietterich, 2012; Feng et al., 2020; Lv et al., 2020), we fix a label (noise) distribution (Assumption 4.1) that permits further analysis of the proposed algorithm. Appendix D experimentally verifies that Assumption 4.1 is satisfied on real-world datasets.
@@
$\operatorname{\mathbb{P}}(S=\underbrace{s}_{\parbox{14.22636pt}{ \tiny\centering\mbox{$=\ \{y\}\ \cup\ z$} \mbox{\text{Partitioning valid in cases $(ii)$\,--\,$(iv)$.}} \@add@centering}}\,,Y=y\mid X=x_{i})=\begin{cases}0&\text{if $s=\operatorname{\mathcal{Y}}$, $s=\emptyset$, $y\notin s$, $y\in z$, or $\tilde{y}\in z$,}\ {\hskip 7.0pt(i)}\\ \frac{1}{2^{l-2}-1}\,p_{1}&\text{if $y=\tilde{y}$, $\tilde{y}_{\operatorname{c}}\in z$, and $s\neq\operatorname{\mathcal{Y}}$,}\ {\hskip 49.57pt(ii)}\\ \frac{1}{2^{l-2}}\,p_{2}&\text{if $y=\tilde{y}$ and $\tilde{y}_{\operatorname{c}}\notin z$,}\ {\hskip 81.125pt(iii)}\\ \frac{1}{2^{l-1}-1}\frac{1}{|s|}\,p_{3}&\text{if $y\neq\tilde{y}$, $\tilde{y}\notin z$, $y\notin z$, and $s\neq\emptyset$,}\ {\hskip 27.6pt(iv)}\\ \end{cases}$
@@
Because $p_{1}+p_{2}>p_{3}$, Assumption 4.1 implies $\operatorname{\mathbb{P}}_{XY}(Y=\tilde{y}\mid X=x_{i})>\operatorname{\mathbb{P}}_{XY}(Y\neq\tilde{y}\mid X=x_{i})$, that is, points that are close in feature space are likely to have similar class labels. In the following, we elaborate on the four cases $(i)$ – $(iv)$.
@@
Theorem 4.3, which establishes risk-consistency, hinges on the following intermediate results in Lemma 4.2. The main difficulty arises in the application of Yager’s rule (4) to produce $\operatorname{\tilde{m}}$ as the focal sets of the $\operatorname{m}_{i}$-s can be arbitrary subsets. The proof of Lemma 4.2 provides a closed-form solution.
@@
To put Theorem 4.3 into context, we recall the concept of risk consistency (Devroye et al., 1996). As Assumption 4.1 fixes a distribution $\operatorname{\mathbb{P}}$, one can quantify the classification risk $\operatorname{\mathcal{R}}(g)$ of a classifier $g$ as defined in Section 2. The Bayes classifier is defined by $g^{*}=\arg\min_{g:\operatorname{\mathcal{X}}\rightarrow\operatorname{\mathcal{Y}}}\operatorname{\mathcal{R}}(g)$; it has the least overall risk. Let $g_{n}$ be the classifier trained by Alg. 1 with $n$ instances. Its empirical risk ^$\hat{\operatorname{\mathcal{R}}}(g_{n})$ can be computed by substituting the expectation with a sample mean. It is risk-consistent if ^$\hat{\operatorname{\mathcal{R}}}(g_{n})\to\operatorname{\mathcal{R}}(g^{*})$ for $n\to\infty$ almost surely. Theorem 4.3 establishes the risk consistency of the proposed classifier.
@@
4.3 Reject Option
To limit the impact of misclassification, our method provides a reject option $\Gamma_{g}$, that is, the algorithm can abstain from individual predictions if unsure (Algorithm 1, Line 12). Recall from (3) that belief describes the lower-bound probability and plausibility the upper-bound probability of the credal set $\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$. Let $\operatorname{\tilde{m}}$ be the resulting bpa as determined by Algorithm 1 and
$\Delta_{\operatorname{\tilde{m}}}:=\operatorname{bel}_{\operatorname{\tilde{m}}}(\{\hat{y}\})-\max_{y\in\tilde{s}\setminus\{\hat{y}\}}\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})\quad\text{with}\quad\hat{y}:=\arg\max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\left\{y\right\})\text{.}$ (5)
$\Delta_{\operatorname{\tilde{m}}}$ satisfies several desirable properties: Lemma 4.4 $(i)$ states that $\Delta_{\operatorname{\tilde{m}}}\leq 0$ if the classification decision has been taken randomly (Algorithm 1, Line 11, second case). Lemma 4.4 $(ii)$ states that, if $\Delta_{\operatorname{\tilde{m}}}>0$, $g(\tilde{x})$ will always be our prediction independent of $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$. Finally, Lemma 4.4 $(iii)$ states that a mass of more than 1/2 is sufficient, but not necessary, for $\Delta_{\operatorname{\tilde{m}}}>0$.
Lemma 4.4.
  1. (i)
  2. (ii)
  3. (iii)
    if $\operatorname{\tilde{m}}(\left\{g(\tilde{x})\right\})>1/2$, then $\Delta_{\operatorname{\tilde{m}}}>0$. The converse does not hold.
Based on these considerations, we define the accept and reject regions of our method accordingly.
  1. 1.
    Accept. When $\Delta_{\operatorname{\tilde{m}}}>0$, we are in the accept region, that is, $\Gamma_{g}(\tilde{x})=\{\hat{y}\}$: The lower-bound probability of the class label $\hat{y}$ is greater than the upper-bound probability of any other class label $y\neq\hat{y}$, that is, ($\operatorname{\mathbb{P}}(\{\hat{y}\})>\operatorname{\mathbb{P}}(\{y\})$ for all $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ and $y\neq\hat{y}$.
  2. 2.
    Reject. When $\Delta_{\operatorname{\tilde{m}}}\leq 0$, we are in the reject region, that is, $\Gamma_{g}(\tilde{x})=\emptyset$: The lower-bound probability of the class label $\hat{y}$ is less than or equal to the upper-bound probability of another class label $y\neq\hat{y}$. There exists $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ and $y\neq\hat{y}$ with ($\operatorname{\mathbb{P}}(\{\hat{y}\})\leq\operatorname{\mathbb{P}}(\{y\})$.
Moreover, Lemma 4.4 $(iii)$ also shows that we reject fewer predictions than the $k$-nearest neighbor reject option by Hellman (1970), which needs always more than 1/2 of all votes to be certain. Our method can accept predictions with less than 1/2 of all probability mass on a single class label, while still satisfying Lemma 4.4 $(ii)$, that is, a prediction remains unchanged independent of $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ if it is accepted.

### v4

@@ 3.2 Dempster-Shafer Theory (DST)
Several methods already leverage DST in supervised learning (Mandler & Schümann, 1988; Denoeux, 1995; Tabassian et al., 2012; Sensoy et al., 2018; Denoeux, 2019; Tong et al., 2021). We consider the nearest neighbor approach by Denoeux (1995) to be most closely related to our approach. Here, basic probability assignments are constructed from the nearest neighbors of an instance. Then, Dempster’s rule is used to combine them into a single bpa. Their analysis is, however, not transferable to our case because they only have singletons or the full label space as focal sets, making set intersections in the combination rule easy to handle. In this sense, we examine a more general setting since we allocate probability mass to arbitrary subsets.
@@ 4 Our Method: DST-PLL
This section introduces our novel partial-label learning method Dst-Pll. Based on the labeling information of an instance’s nearest neighbors, we construct basic probability assignments within Dempster-Shafer theory. These bpas inform the prediction and rejection decisions as discussed in Section 4.1 and Section 4.2, respectively. Regarding the reject option, we propose a novel variation of the confidence-based rejection strategy: The confidence threshold is adaptively selected on a per-instance basis dependent on the amount of incorrect label noise. The more noise from incorrect labels there is, the more confident the model needs to be to accept a prediction.
Algorithm 1 outlines Dst-Pll, which we summarize in the following. We denote by $\operatorname{NN}_{k}(\tilde{x})\subseteq\operatorname{\mathcal{X}}\times 2^{\operatorname{\mathcal{Y}}}$ the set of the $k$-nearest neighbors of instance $\tilde{x}$ with their associated candidate labels. To predict the class label of an instance $\tilde{x}$ (Line 11), the algorithm first transforms information from $\tilde{x}$’s neighbors $\operatorname{NN}_{k}(\tilde{x})$ into bpas $\operatorname{m}_{i}$ (Lines 3–7), collects these into evidence set $\mathcal{E}$ (Line 8), and combines the bpas into $\operatorname{\tilde{m}}$ using Yager’s rule (Line 10; Yager 1987a; b). Section 4.1 elaborates on these steps. Section 4.2 elaborates on how we extract our reject option from $\operatorname{\tilde{m}}$ (Line 12). Section 4.3 shows that the resulting classification rule is risk-consistent. We analyze our algorithm’s runtime in Section 4.4.
@@
4.1 Making Predictions
Basic probability assignments. Following the standard assumption that neighboring instances in feature space are also close in label space, we combine the evidence from the $k$-nearest neighbors $(x_{i},s_{i})\in\operatorname{NN}_{k}(\tilde{x})$ of a given instance $\tilde{x}\in\operatorname{\mathcal{X}}$ with its candidate labels $\tilde{s}\subseteq\operatorname{\mathcal{Y}}$ ($\tilde{s}=\operatorname{\mathcal{Y}}$ if $\tilde{x}$ is a test instance).
When looking at a neighboring instance $(x_{i},s_{i})\in\operatorname{NN}_{k}(\tilde{x})$, there are generally two cases: either (i) $(x_{i},s_{i})$ provides information about the correct label of $\tilde{x}$ or (ii) $(x_{i},s_{i})$ is irrelevant for finding the correct label of $\tilde{x}$. To address (i), we allocate probability mass on the candidates $s_{i}$ of neighbor $x_{i}$. To address (ii), we allocate probability mass on the full label space $\operatorname{\mathcal{Y}}$ indicating uncertainty about the correct label of $\tilde{x}$.
More formally, for fixed $i\in[k]$, the candidate labels $s_{i}$ do not provide any valuable information if they support all ($\tilde{s}\subseteq s_{i}$) or none ($\tilde{s}\cap s_{i}=\emptyset$) of the labels in $\tilde{s}$; we use a bpa of $\operatorname{m}_{i}(\tilde{s})=1$ (Line 4). We set $\operatorname{m}_{i}(A)=1/2$ if $A=\tilde{s}$ or $A=\tilde{s}\cap s_{i}$, else $\operatorname{m}_{i}(A)=0$ (Line 6), where $1/2$ equally weights evidence. We later elaborate further on this choice and demonstrate the application of the proposed classification rule in Example 4.1. Note that we make the common assumption that the true label of instance $\tilde{x}$ is always in $\tilde{s}$ (Cour et al., 2011; Liu & Dietterich, 2012; Lv et al., 2020; Ni et al., 2021). While our definition of the $\operatorname{m}_{i}$-s is similar to (Denoeux, 1995), we target a more general setting as our focal sets can be arbitrary subsets instead of only singletons or the full label set.
@@ 4.1 Making Predictions
Evidence combination. Given the set =$\mathcal{E}=\{\operatorname{m}_{i}\mid i\in[k]\}$, we combine all $\operatorname{m}_{i}$-s using Yager’s rule (Yager, 1987a; b). Dempster’s original rule (Dempster, 1967) enforces $\operatorname{\tilde{m}}(\emptyset)=0$ by normalization, which is criticized for its unintuitive results when facing high conflict (Zadeh, 1984). Instead, Yager’s rule first collects overlapping evidence in $\operatorname{q}:2^{\operatorname{\mathcal{Y}}}\rightarrow[0,1]$ and creates a valid bpa $\operatorname{\tilde{m}}:2^{\operatorname{\mathcal{Y}}}\rightarrow[0,1]$ by
@@
Classification rule. After the combination into $\operatorname{\tilde{m}}$ by (4), we extract a prediction $g(\tilde{x})$ for instance $\tilde{x}$ (Line 11). We predict the class label with the highest probability mass $\arg\max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\{y\})$ if any has non-zero mass, or else randomly pick from the subset with the most mass $\arg\max_{A\subseteq\tilde{s}}\,\operatorname{\tilde{m}}(A)$. When tied, we use the subset with the smallest cardinality. In the following, we present an example of our classification rule.
Example 4.1 (Classification rule).
Let $k=3$, $\tilde{x}$ an unseen test instance, $(x_{i},s_{i})\in\operatorname{NN}_{k}(\tilde{x})$, $\operatorname{\mathcal{Y}}=\{1,2,3\}$, $s_{1}=\{1\}$, $s_{2}=\{1,2\}$, and $s_{3}=\{1,3\}$. Then, $\operatorname{m}_{1}(\{1\})=\operatorname{m}_{1}(\operatorname{\mathcal{Y}})=1/2$, $\operatorname{m}_{2}(\{1,2\})=\operatorname{m}_{2}(\operatorname{\mathcal{Y}})=1/2$, and $\operatorname{m}_{3}(\{1,3\})=\operatorname{m}_{3}(\operatorname{\mathcal{Y}})=1/2$. All other subsets receive a mass of zero. Using Yager’s combination rule, we obtain $\operatorname{\tilde{m}}(\{1\})=5/8$, $\operatorname{\tilde{m}}(\{1,2\})=\operatorname{\tilde{m}}(\{1,3\})=\operatorname{\tilde{m}}(\operatorname{\mathcal{Y}})=1/8$, and $\operatorname{\tilde{m}}(A)=0$ for the remaining $A\subseteq\operatorname{\mathcal{Y}}$. Therefore, we predict label $1$.
4.2 Proposed Reject Option
To limit the impact of misclassification, our method provides a reject option $\Gamma_{g}$, that is, the algorithm can abstain from individual predictions if unsure (Algorithm 1, Line 12). Our formulation builds on the confidence-based rejection strategy (Section 2.2), that is, $\Delta:=\operatorname{conf}(g)-\theta$ with $\operatorname{conf}(g)\in[0,1]$ being the model’s confidence and $\theta\in[0,1]$ the confidence threshold. We adapt this setting to the PLL context by changing the confidence threshold $\theta_{\operatorname{\tilde{m}}}$ based on the amount of noise present.
Recall from (3) that the belief and plausibility regarding $\operatorname{\tilde{m}}$ act as a lower and upper bound of the probability mass, respectively. The intuition of our reject option is as follows. If the lower bound (belief) on the probability mass of our predicted label exceeds the maximal upper bound (plausibility) on the probability mass regarding any other label, we can safely make the prediction.
In other words, if there is a class label different from the predicted one that is quite plausible (high $\max_{y\in\tilde{s}\setminus\{\hat{y}\}}\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})$), we require a high belief mass in order to be certain about the prediction, that is, the belief must satisfy $\operatorname{bel}_{\operatorname{\tilde{m}}}(\{\hat{y}\})>\max_{y\in\tilde{s}\setminus\{\hat{y}\}}\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})$. If, instead, there is no other plausible candidate label, we can be sure of our prediction with less belief mass.
We formalize this intuition in the following. Let $\operatorname{\tilde{m}}$ be the resulting bpa as determined by Algorithm 1 and
$\Delta_{\operatorname{\tilde{m}}}:=\underbrace{\operatorname{bel}_{\operatorname{\tilde{m}}}(\{\hat{y}\})}_{(=\operatorname{conf}(g))}-\underbrace{\max_{y\in\tilde{s}\setminus\{\hat{y}\}}\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})}_{(=\theta_{\operatorname{\tilde{m}}})}\quad\text{with}\quad\hat{y}:=\arg\max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\left\{y\right\})\text{.}$ (5)
In other words, we instantiate the model’s confidence with the model’s belief mass $\operatorname{conf}(g)=\operatorname{bel}_{\operatorname{\tilde{m}}}(\{\hat{y}\})$ of the predicted instance $\hat{y}$ and the confidence threshold $\theta_{\operatorname{\tilde{m}}}$ based on the amount of noise regarding other labels $y\neq\hat{y}$, that is, $\theta_{\operatorname{\tilde{m}}}=\max_{y\in\tilde{s}\setminus\{\hat{y}\}}\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})$. Note that the dependence on $\operatorname{\tilde{m}}$ allows for setting the threshold adaptively.
(5) satisfies several desirable properties, which we collect in Theorem 4.2 and elaborate on in the following.
Theorem 4.2.
  1. (i)
  2. (ii)
  3. (iii)
    if $\operatorname{\tilde{m}}(\left\{g(\tilde{x})\right\})>1/2$, then $\Delta_{\operatorname{\tilde{m}}}>0$. The converse of (iii) does not hold.
Based on these considerations, we define the accept and reject regions of our method as follows.
  1. 1.
    Accept. When $\Delta_{\operatorname{\tilde{m}}}>0$, we are in the accept region and $\Gamma_{g}(\tilde{x})=\{\hat{y}\}$: The lower-bound probability of the class label $\hat{y}$ is greater than the upper-bound probability of any other class label $y\neq\hat{y}$, that is, ($\operatorname{\mathbb{P}}(\{\hat{y}\})>\operatorname{\mathbb{P}}(\{y\})$ for all $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ and $y\neq\hat{y}$.
  2. 2.
    Reject. When $\Delta_{\operatorname{\tilde{m}}}\leq 0$, we are in the reject region and $\Gamma_{g}(\tilde{x})=\emptyset$: The lower-bound probability of the class label $\hat{y}$ is less than or equal to the upper-bound probability of another class label $y\neq\hat{y}$. There exists $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$ and $y\neq\hat{y}$ with ($\operatorname{\mathbb{P}}(\{\hat{y}\})\leq\operatorname{\mathbb{P}}(\{y\})$.
We remark that Theorem 4.2 $(iii)$ implies that Dst-Pll rejects fewer predictions than the $k$-nearest neighbor reject option by Hellman (1970), which requires more than 1/2 of all votes to be certain. Our method can accept predictions with less than 1/2 of all probability mass on a single class label, while still satisfying Theorem 4.2 $(ii)$, that is, if a prediction is accepted, the decision remains unchanged independent of $\operatorname{\mathbb{P}}\in\operatorname{\mathcal{C}}_{\operatorname{\tilde{m}}}(\operatorname{\mathcal{Y}},2^{\operatorname{\mathcal{Y}}})$. It follows that having fewer rejections does not come at the expense of an increase in the expected error. In the following, we provide an example of the proposed reject option.
Example 4.3 (Reject option).
Assume the setting and result of Example 4.1: $\operatorname{\tilde{m}}(\{1\})=5/8$, $\operatorname{\tilde{m}}(\{1,2\})=\operatorname{\tilde{m}}(\{1,3\})=\operatorname{\tilde{m}}(\operatorname{\mathcal{Y}})=1/8$, and $\operatorname{\tilde{m}}(A)=0$ for the remaining $A\subseteq\operatorname{\mathcal{Y}}$. Our prediction is $\hat{y}=\arg\max_{y\in\tilde{s}}\,\operatorname{\tilde{m}}(\left\{y\right\})=1$ with $\tilde{s}=\operatorname{\mathcal{Y}}$. Hence, $\Delta_{\operatorname{\tilde{m}}}=\operatorname{bel}_{\operatorname{\tilde{m}}}(\{\hat{y}\})-\max_{y\in\tilde{s}\setminus\{\hat{y}\}}\operatorname{pl}_{\operatorname{\tilde{m}}}(\{y\})=\operatorname{bel}_{\operatorname{\tilde{m}}}(\{1\})-\operatorname{pl}_{\operatorname{\tilde{m}}}(\{2\})=5/8-2/8=3/8$. Since $\Delta_{\operatorname{\tilde{m}}}=3/8>0$, we accept the prediction $\hat{y}=1$.
4.3 Consistency
@@
Our classification rule yields a risk-consistent classifier, which we demonstrate in the following. As is common in the literature (Cour et al., 2011; Liu & Dietterich, 2012; Feng et al., 2020; Lv et al., 2020) and required to obtain statistical guarantees in the PLL setting, we fix a label (noise) distribution (Assumption 4.4) that permits further analysis of the proposed algorithm. Appendix D experimentally verifies that Assumption 4.4 is satisfied on real-world datasets.
@@
$\operatorname{\mathbb{P}}(S=\underbrace{s}_{\parbox{14.22636pt}{ \tiny\centering\mbox{$=\ \{y\}\ \cup\ z$} \mbox{\text{Partitioning valid in cases $(ii)$\,--\,$(iv)$.}} \@add@centering}}\,,Y=y\mid X=x_{i})=\begin{cases}0&\text{if $s=\operatorname{\mathcal{Y}}$, $s=\emptyset$, $y\notin s$, $y\in z$, or $\tilde{y}\in z$,}\ {\hskip 6.3pt(i)}\\ \frac{1}{2^{l-2}-1}\,p_{1}&\text{if $y=\tilde{y}$, $\tilde{y}_{\operatorname{c}}\in z$, and $s\neq\operatorname{\mathcal{Y}}$,}\ {\hskip 50.0pt(ii)}\\ \frac{1}{2^{l-2}}\,p_{2}&\text{if $y=\tilde{y}$ and $\tilde{y}_{\operatorname{c}}\notin z$,}\ {\hskip 83.0pt(iii)}\\ \frac{1}{2^{l-1}-1}\frac{1}{|s|}\,p_{3}&\text{if $y\neq\tilde{y}$, $\tilde{y}\notin z$, $y\notin z$, and $s\neq\emptyset$,}\ {\hskip 27.3pt(iv)}\\ \end{cases}$
@@
Because $p_{1}+p_{2}>p_{3}$, Assumption 4.4 implies $\operatorname{\mathbb{P}}_{XY}(Y=\tilde{y}\mid X=x_{i})>\operatorname{\mathbb{P}}_{XY}(Y\neq\tilde{y}\mid X=x_{i})$, that is, points that are close in feature space are likely to have similar class labels. We note that this assumption is related to the ambiguity degree condition by Cour et al. (2011) as both make sure that the noise labels do not overwhelm the PLL algorithm. Assumption 4.4 enforces this as $p_{1}\geq p_{2}\geq p_{3}>0$. Having all cases spelled out as in Assumption 4.4 benefits the proof of Lemma 4.5.
In the following, we elaborate on the four cases $(i)$ – $(iv)$.
@@
Theorem 4.6, which establishes risk-consistency, hinges on the following intermediate results in Lemma 4.5. The fact that the $\operatorname{m}_{i}$-s can have arbitrary focal sets renders the direct application of Yager’s rule (4) in our analysis challenging. The proof of Lemma 4.5 mediates this issue by providing a closed-form solution for the resulting bpa $\operatorname{\tilde{m}}$.
@@
To put Theorem 4.6 into context, we recall the concept of risk consistency (Devroye et al., 1996). The Bayes classifier is defined by $g^{*}=\arg\min_{g:\operatorname{\mathcal{X}}\rightarrow\operatorname{\mathcal{Y}}}\operatorname{\mathcal{R}}(g)$; it has the least overall risk. Let $g_{n}$ be the classifier trained by Alg. 1 with $n$ instances. Its empirical risk ^$\hat{\operatorname{\mathcal{R}}}(g_{n})$ can be computed by substituting the expectation with a sample mean. It is risk-consistent if ^$\hat{\operatorname{\mathcal{R}}}(g_{n})\to\operatorname{\mathcal{R}}(g^{*})$ for $n\to\infty$ almost surely. Theorem 4.6 establishes the risk consistency of the proposed classifier.